Nitrous acid, HNO2(aq), has a Ka of 4.5 x 10-4 at 25 °C. What is the pH of a solution that contains 1.0 M KNO2(aq) and 0.50 M HNO2(aq).
Answer #1

(acid) - [HNO2) = 0.50m (salt) = [KNO2) = 1.0M Ka= 4.5X10 4 pre = -log ka = -log(4.5 x104) pH = pka + log (Salt PH = -log(4.5